OpenStudy (anonymous):

solve absolute value of x-10=22

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

\[\left| x-10 \right|=22\]

OpenStudy (zepp):

If |x-10| = 22 it implies that \(x-10\) gives two answers, 22 and -22 So let's write out our first answer x-10 = 22 and solve for x Second answer... x-10 = -22 and solve for x

OpenStudy (anonymous):

how..?

OpenStudy (anonymous):

@nasavivek x=-12,x=12/or x=-12,x=32